OpenStudy (anonymous):

A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected?

OpenStudy (rulnick):

20^n=10^8 nlog20=8 n=8/log20=6.15 (approx) So it will take about 6.15 times 5 minutes, or about 30.74 minutes. Since the problem says "every five minutes" you may need to round to 35 minutes.